# On the first-order genus of wreath products and their central extensions

### 16 Mar 2026 v1

@@ 3.2 Interpretability of $\mathbb{Z}$} in $G$
In particular, the formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})=Prod_{A,2}(x_{1},x_{2},\bar{y},\bar{z})$ defines in $G$ the $\mathbb{Z}$}-exponentiation on $A$, that is for any $a,d\neq 1\in A$ and any $\gamma\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
@@ 4.1 Definability of $\mathbb{Z}$}-exponentiations in $G$
Moreover, $w=u^{k}$ is the only element in $N$ such that for every $a\in A$ the equation
@@
Since $N$ is a free ${\mathbb{Z}}A$-module, one gets $w^{-1}u^{k}=1.$ Indeed, let $w^{-1}u^{k}=\Pi_{i}b_{i}^{P_{i}}$ and $v_{1}v^{-1}=\Pi_{i}b_{i}^{Q_{i}}$ for some $P_{i},Q_{i}\in{\mathbb{Z}}A$. Then for every $i$ $P_{i}=Q_{i}(a-1)$. It follows that a1a-1 divides $Q_{i}$ for every $a\in A$. In particular, every polynomial $a_{1}^{\ell}-1,\ell>0$ divides every $P_{i}$. Hence, $P_{i}=0$ for all $i$, so $w^{-1}u^{k}=1$, as claimed. ∎
Corollary 4.4.
@@
4.2 Interpretation of the ${\mathbb{Z}}A$-action on $N$ in $G$
In this section, we show that the standard action of ${\mathbb{Z}}A$ on $N$ is interpretable in $G$ with parameters $a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}$. We fix an element $a=a_{1}\in A$ and use the interpretation ${\mathbb{Z}}_{a}$ of the ring of integers $\mathbb{Z}$} in $G$.
@@ 4.2 Interpretation of the ${\mathbb{Z}}A$-action on $N$ in $G$
Denote by $(N)^{I_{\bar{\alpha}}}$ the submodule of the ${\mathbb{Z}}A$-module $N$ generated by $N^{I_{\bar{\alpha}}}$. Observe, that every element $u\in N^{I_{\bar{\alpha}}}$, as well as in $(N)^{I_{\bar{\alpha}}}$ admits a decomposition of the type $u=\Pi_{i}b_{i}^{U_{i}}$, where $U_{i}\in I_{\bar{\alpha}}$. As an abelian group, $(N)^{I_{\bar{\alpha}}}$ is generated by the set $\{g^{Q}\mid g\in N,Q\in I_{\bar{\alpha}}\}$, hence by the set $\{g^{a_{i}-\alpha_{i}}\mid g\in N,i=1,\ldots,m\}$. The result below shows that the subgroup $(N)^{I_{\bar{\alpha}}}$ is definable in $G$ uniformly in $a_{1},\ldots,a_{m}$ and $(\alpha_{1},\ldots,\alpha_{m})$. Of course, integers are not elements of $G$, so instead of $(\alpha_{1},\ldots,\alpha_{m})$ we use their interpretations $a^{\alpha_{1}},\ldots,a^{\alpha_{m}}$ in ${\mathbb{Z}}_{a}$.
@@
There is a formula $\phi(x,\bar{y},\bar{z})$ of the group language such that for any basis $(a_{1},\ldots,a_{m})$ of $A$ and any tuple $(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the formula
@@
Since $N$ is free with basis $b_{1},\ldots,b_{n}$ it follows from (17) that for $i=1,\ldots,n$
@@
Now we can prove that in the group $G$ the action of ${\mathbb{Z}}A$ on $N$ is definable. Again, Laurent polynomials $Q\in{\mathbb{Z}}A$ are not elements of $G$, so we first represent them by their codes $k=\nu^{-1}(Q)$, and then represent the number $k$ as $a^{k}$ in the interpretation ${\mathbb{Z}}_{a}$ of $\mathbb{Z}$} in $G$. As above, we may choose $a=a_{1}$. The result is stated as follows.
@@
This defines the action of the ring of ordinary polynomials ${\mathbb{Z}}[\bar{a}]$ on $N$.
The last step is to get a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ that defines the action of the ring of the Laurent polynomials ${\mathbb{Z}}A$ on $N$. Let $Q\in{\mathbb{Z}}A$. Then $Q=\frac{P}{M}$, where $P,M\in{\mathbb{Z}}[\bar{a}]$ and $M$ is a monomial. In this case, $g^{Q}=h$ if and only if $g^{P}=h^{M}$, and the exponentiation here by the polynomials $P$ and $M$, so the formula $Act_{{\mathbb{Z}}[\bar{a}]}(x_{1},x_{2},x_{3},y_{1})$ does it for us.
@@
There is a formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ of group language such that for any $u\in N$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
@@
There is a formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ of group language such that for any $g\in G$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of $N$, any tuple $\bar{\gamma}=(\gamma_{1},\ldots,\gamma_{m})\in{\mathbb{Z}}^{m}$, and any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
@@
By Corollary 4.10 for $u\in N$ and $P_{1},\ldots,P_{n}\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ the condition $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ from (21) is described by the formula $\Lambda_{N}$. Using this one can easily write a formula that states that the element $u$ such that $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ is unique in $N$, which means that $N$ is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
@@ 5 Bi-interpretability of $G$ and $\mathbb{Z}$}
Fix a tuple of parameters $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively. In this section we prove that $G$ and $\mathbb{Z}$} are bi-interpretable with parameters $\bar{p}$ in $G$.
@@
    Let $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of $A$ and $B$ respectively. Then the group $G$ and the ring $\mathbb{Z}$} are bi-interpretable with each other through the interpretations $G\simeq\Delta({\mathbb{Z}})$ and ${\mathbb{Z}}\simeq\Gamma(G,\bar{p})$.
@@ 6 Condensed center-by-metabelian groups
We begin by recalling some definitions from [13]. For an abstract set $I$, we denote by $2^{I}$ the space of all subsets of $I$ endowed with the product topology (equivalently, the topology of pointwise convergence of the indicator functions). If $G$ is a group, we denote by $\mathcal{N}(G)$ the set of all normal subgroups of $G$ endowed with the subspace topology induced by the inclusion $\mathcal{N}(G)\subseteq 2^{G}$. Thus, the base of neighborhoods of $N\in\mathcal{N}(G)$ consists of the sets
@@ 6.2 The main construction
    The center $Z(G)$ is a free group in the variety of abelian groups of exponent $2$, and $\{c_{k}\mid k\in\mathbb{Z}\}$ is a basis in $Z(G)$.
@@
Let $N$ be the free nilpotent group of rank $2$ with basis $\{\tilde{a}_{i},\tilde{b}_{i}\mid i\in\mathbb{Z}\}$. It is well-known that $[N,N]=Z(N)$ is a free abelian group freely generated by the basic commutators $\tilde{a}_{i},\tilde{a}_{j}]$, $\tilde{b}_{i},\tilde{b}_{j}]$ for $j<j$, $i,j\in\mathbb{Z}$, and $\tilde{a}_{i},\tilde{b}_{j}]$ for $i,j\in\mathbb{Z}$ (see [29, Ch. 3, Sec. 1] and references therein). Let $M\leq[N,N]$ be the subgroup generated by the elements
@@
for all $i,j,\ell\in\mathbb{Z}$. It is straightforward to verify that $[N,N]/M$ is a free group in the variety of abelian groups of exponent $2$, freely generated by elements $c_{k}=[\tilde{a}_{0},\tilde{b}_{k}]M$, $k\in\mathbb{Z}$.
Since $M$ is central in $N$, we can consider the quotient group $N/M$. We denote the images of $\tilde{a}_{i}$ and $\tilde{b}_{i}$ ($i,j\in\mathbb{Z}$) in $N/M$ by $a_{i}$ and $b_{j}$, respectively. Our construction yields the presentation
@@
where $A$ is given by (23) and the indices $i$, $j$ independently range in $\mathbb{Z}$.
@@
To prove injectivity, let $S$ and $T$ be two distinct subsets of $\mathbb{Z}$. Without loss of generality, we can assume that there exists $i\in S\setminus T$. By Lemma 6.7, the relation $a,b_{0}^{t^{i}}]=1$ holds in $G_{T}$. On the other hand, the element $a,b_{0}^{t^{i}}]=c_{i}$ generates $Z(G_{S})\cong\mathbb{Z}_{2}$ by Lemma 6.6. In particular, $a,b_{0}^{t^{i}}]\neq 1$ in $G_{S}$. By the definition of marked groups, this implies that $(G_{S},\{a,b_{0},t\})$ and $(G_{T},\{a,b_{0},t\})$ represent distinct elements of $\mathcal{G}$.
@@
Let $\mathcal{C}=f(2^{\mathbb{Z}})$, where $f$ is the map from Lemma 6.9. Note that $\mathcal{C}$ consists of groups of the form (1) by Lemma 6.6. Since $2^{\mathbb{Z}}$ is compact and $f$ is continuous, $\mathcal{C}$ is compact; in particular, $\mathcal{C}$ is a closed subset of $\mathcal{G}$. Note also that $\mathcal{C}$ has no isolated points being the image of a space without isolated points under a continuous and injective map.

### 18 Mar 2026 v2

@@ 3.2 Interpretability of $\mathbb{Z}$} in $G$
In particular, the formula $Exp_{A}(x_{1},x_{2},y_{1},z_{1})=Prod_{A,2}(x_{1},x_{2},\bar{y},\bar{z})$ defines in $G$ the $\mathbb{Z}$}-exponentiation on AA, that is for any $a,d\neq 1\in A$ and any $\gamma\in{\mathbb{Z}}$, for any $g\in G$ the following holds:
@@ 4.1 Definability of $\mathbb{Z}$}-exponentiations in $G$
Moreover, $w=u^{k}$ is the only element in NN such that for every $a\in A$ the equation
@@
Since NN is a free ${\mathbb{Z}}A$-module, one gets $w^{-1}u^{k}=1.$ Indeed, let $w^{-1}u^{k}=\Pi_{i}b_{i}^{P_{i}}$ and $v_{1}v^{-1}=\Pi_{i}b_{i}^{Q_{i}}$ for some $P_{i},Q_{i}\in{\mathbb{Z}}A$. Then for every $i$ $P_{i}=Q_{i}(a-1)$. It follows that a1a-1 divides $Q_{i}$ for every $a\in A$. In particular, every polynomial $a_{1}^{\ell}-1,\ell>0$ divides every $P_{i}$. Hence, $P_{i}=0$ for all $i$, so $w^{-1}u^{k}=1$, as claimed. ∎
Corollary $4.$
@@
4.2 Interpretation of the ${\mathbb{Z}}A$-action on NN in $G$
In this section, we show that the standard action of ${\mathbb{Z}}A$ on NN is interpretable in $G$ with parameters $a_{1},\ldots,a_{m},b_{1},\ldots,b_{n}$. We fix an element $a=a_{1}\in A$ and use the interpretation ${\mathbb{Z}}_{a}$ of the ring of integers $\mathbb{Z}$} in $G$.
@@ 4.2 Interpretation of the ${\mathbb{Z}}A$-action on NN in $G$
Denote by $(N)^{I_{\bar{\alpha}}}$ the submodule of the ${\mathbb{Z}}A$-module NN generated by $N^{I_{\bar{\alpha}}}$. Observe, that every element $u\in N^{I_{\bar{\alpha}}}$, as well as in $(N)^{I_{\bar{\alpha}}}$ admits a decomposition of the type $u=\Pi_{i}b_{i}^{U_{i}}$, where $U_{i}\in I_{\bar{\alpha}}$. As an abelian group, $(N)^{I_{\bar{\alpha}}}$ is generated by the set $\{g^{Q}\mid g\in N,Q\in I_{\bar{\alpha}}\}$, hence by the set $\{g^{a_{i}-\alpha_{i}}\mid g\in N,i=1,\ldots,m\}$. The result below shows that the subgroup $(N)^{I_{\bar{\alpha}}}$ is definable in $G$ uniformly in $a_{1},\ldots,a_{m}$ and $(\alpha_{1},\ldots,\alpha_{m})$. Of course, integers are not elements of $G$, so instead of $(\alpha_{1},\ldots,\alpha_{m})$ we use their interpretations $a^{\alpha_{1}},\ldots,a^{\alpha_{m}}$ in ${\mathbb{Z}}_{a}$.
@@
There is a formula $\phi(x,\bar{y},\bar{z})$ of the group language such that for any basis $(a_{1},\ldots,a_{m})$ of AA and any tuple $(\alpha_{1},\ldots,\alpha_{m})\in{\mathbb{Z}}^{m}$ the formula
@@
Since NN is free with basis $b_{1},\ldots,b_{n}$ it follows from (17) that for $i=1,\ldots,n$
@@
Now we can prove that in the group $G$ the action of ${\mathbb{Z}}A$ on NN is definable. Again, Laurent polynomials $Q\in{\mathbb{Z}}A$ are not elements of $G$, so we first represent them by their codes $k=\nu^{-1}(Q)$, and then represent the number $k$ as $a^{k}$ in the interpretation ${\mathbb{Z}}_{a}$ of $\mathbb{Z}$} in $G$. As above, we may choose $a=a_{1}$. The result is stated as follows.
@@
This defines the action of the ring of ordinary polynomials ${\mathbb{Z}}[\bar{a}]$ on NN.
The last step is to get a formula $Act_{{\mathbb{Z}}A}(x_{1},x_{2},x_{3},y_{1})$ that defines the action of the ring of the Laurent polynomials ${\mathbb{Z}}A$ on NN. Let $Q\in{\mathbb{Z}}A$. Then $Q=\frac{P}{M}$, where $P,M\in{\mathbb{Z}}[\bar{a}]$ and $M$ is a monomial. In this case, $g^{Q}=h$ if and only if $g^{P}=h^{M}$, and the exponentiation here by the polynomials $P$ and $M$, so the formula $Act_{{\mathbb{Z}}[\bar{a}]}(x_{1},x_{2},x_{3},y_{1})$ does it for us.
@@
There is a formula $\Lambda_{N}(x,\bar{z},\bar{t},\bar{s})$ of group language such that for any $u\in N$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of NN, any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
@@
There is a formula $\Lambda_{G}(x,\bar{x},\bar{y},\bar{z},\bar{t},\bar{s})$ of group language such that for any $g\in G$, any top basis $\bar{c}=(c_{1},\ldots,c_{m})$ of $G$, any tuple $\bar{u}=(u_{1},\ldots,u_{n})$ of elements of NN, any tuple $\bar{\gamma}=(\gamma_{1},\ldots,\gamma_{m})\in{\mathbb{Z}}^{m}$, and any tuple $\bar{P}=(P_{1},\ldots,P_{n})\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ one has
@@
By Corollary 4.10 for $u\in N$ and $P_{1},\ldots,P_{n}\in{\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$ the condition $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ from (21) is described by the formula $\Lambda_{N}$. Using this one can easily write a formula that states that the element $u$ such that $u=u_{1}^{P_{1}}\ldots u_{n}^{P_{n}}$ is unique in NN, which means that NN is a free ${\mathbb{Z}}[c_{1}^{\pm 1},\ldots,c_{m}^{\pm 1}]$-module with basis $u_{1},\ldots,u_{n}$.
@@ 5 Bi-interpretability of $G$ and $\mathbb{Z}$}
Fix a tuple of parameters $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of AA and $B$ respectively. In this section we prove that $G$ and $\mathbb{Z}$} are bi-interpretable with parameters $\bar{p}$ in $G$.
@@
    Let $p=(a_{1},\ldots,a_{m},b_{1},\ldots,b_{n})$ in $G$, where $\{a_{1},\ldots,a_{m}\}$ and $\{b_{1},\ldots,b_{n}\}$ are bases of AA and $B$ respectively. Then the group $G$ and the ring $\mathbb{Z}$} are bi-interpretable with each other through the interpretations $G\simeq\Delta({\mathbb{Z}})$ and ${\mathbb{Z}}\simeq\Gamma(G,\bar{p})$.
@@ 6 Condensed center-by-metabelian groups
We begin by recalling some definitions from [13]. For an abstract set II, we denote by $2^{I}$ the space of all subsets of II endowed with the product topology (equivalently, the topology of pointwise convergence of the indicator functions). If $G$ is a group, we denote by $\mathcal{N}(G)$ the set of all normal subgroups of $G$ endowed with the subspace topology induced by the inclusion $\mathcal{N}(G)\subseteq 2^{G}$. Thus, the base of neighborhoods of $N\in\mathcal{N}(G)$ consists of the sets
@@ 6.2 The main construction
    The center $Z(G)$ is a free group in the variety of abelian groups of exponent 22, and $\{c_{k}\mid k\in\mathbb{Z}\}$ is a basis in $Z(G)$.
@@
Let NN be the free nilpotent group of rank 22 with basis $\{\tilde{a}_{i},\tilde{b}_{i}\mid i\in\mathbb{Z}\}$. It is well-known that $[N,N]=Z(N)$ is a free abelian group freely generated by the basic commutators $\tilde{a}_{i},\tilde{a}_{j}]$, $\tilde{b}_{i},\tilde{b}_{j}]$ for $j<j$, $i,j\in\mathbb{Z}$, and $\tilde{a}_{i},\tilde{b}_{j}]$ for $i,j\in\mathbb{Z}$ (see [29, Ch. 3, Sec. 1] and references therein). Let $M\leq[N,N]$ be the subgroup generated by the elements
@@
for all $i,j,\ell\in\mathbb{Z}$. It is straightforward to verify that $[N,N]/M$ is a free group in the variety of abelian groups of exponent 22, freely generated by elements $c_{k}=[\tilde{a}_{0},\tilde{b}_{k}]M$, $k\in\mathbb{Z}$.
Since $M$ is central in NN, we can consider the quotient group $N/M$. We denote the images of $\tilde{a}_{i}$ and $\tilde{b}_{i}$ ($i,j\in\mathbb{Z}$) in $N/M$ by $a_{i}$ and $b_{j}$, respectively. Our construction yields the presentation
@@
where AA is given by (23) and the indices $i$, $j$ independently range in $\mathbb{Z}$.
@@
To prove injectivity, let $S$ and $T$ be two distinct subsets of $\mathbb{Z}$. Without loss of generality, we can assume that there exists $i\in S\setminus T$. By Lemma 6.7, the relation $a,b_{0}^{t^{i}}]=1$ holds in $G_{T}$. On the other hand, the element $a,b_{0}^{t^{i}}]=c_{i}$ generates $Z(G_{S})\cong\mathbb{Z}_{2}$ by Lemma $6.$ In particular, $a,b_{0}^{t^{i}}]\neq 1$ in $G_{S}$. By the definition of marked groups, this implies that $(G_{S},\{a,b_{0},t\})$ and $(G_{T},\{a,b_{0},t\})$ represent distinct elements of $\mathcal{G}$.
@@
Let $\mathcal{C}=f(2^{\mathbb{Z}})$, where $f$ is the map from Lemma 6.9. Note that $\mathcal{C}$ consists of groups of the form (1) by Lemma $6.$ Since $2^{\mathbb{Z}}$ is compact and $f$ is continuous, $\mathcal{C}$ is compact; in particular, $\mathcal{C}$ is a closed subset of $\mathcal{G}$. Note also that $\mathcal{C}$ has no isolated points being the image of a space without isolated points under a continuous and injective map.